OpenStudy (anonymous):

Identify the y and x-intercepts of g(x). f(x) = x^2 -4 g(x) = f(x)-5

OpenStudy (anonymous):

To find x - intercept, let y = 0 and solve for x, for y - intercept, let x = 0 and solve for y. g(x) = (x^2-4)-5 since f(x) = x^2-4

OpenStudy (anonymous):

Im so lost..I been struggling over this all yesterday and decided to give it a break.

OpenStudy (anonymous):

That's ok, can you show me what you tried?

OpenStudy (anonymous):

I know that the x intercept of F(x)=x^2-4 is -2 and 2. And that the y intercept is -4. But i don't know where to go from there.

OpenStudy (anonymous):

@iambatman ?

OpenStudy (anonymous):

You're just plugging in f(x) in g(x) and essentially doing the same thing

OpenStudy (anonymous):

You will see g(x) becomes (x^2-4)-5 and you can find the x and y intercepts as before

OpenStudy (anonymous):

So 3 and -3 and y intercept is -9?

OpenStudy (anonymous):

Looks good to me!